OpenStudy (beccaboo333):

Part 1: Write the equation of the line that passes through the points (3, 4) and (-3, -8) in point-slope form. Part 2: Using complete sentences, explain whether or not it matters which point is used in the final answer. Also explain why you chose the point you did.

OpenStudy (plainntall):

Part 1. What is point slope form?

OpenStudy (beccaboo333):

i have no idea

OpenStudy (plainntall):

Is slope the change in y divided by the change in x or rise or run?

OpenStudy (beccaboo333):

honestly i dont know

OpenStudy (plainntall):

It is. Point slope for is y-y1=m(x-x1) where m is the slope

OpenStudy (anonymous):

point slope is \[y-y _{1}=m(x-x _{1})\]

OpenStudy (plainntall):

so if m=(y2-y1)/(x2-x1) where (x2,y2)= (3, 4) and (x1,y1)=(-3,-8) what is m?

OpenStudy (anonymous):

so the slope would be y2-y1/x2-x1 , right?

OpenStudy (anonymous):

point = (x,y) an ordered pair when you know the slope (m) you can create the equation of the line y = mx+b where b is the y intersect at x=0

OpenStudy (anonymous):

so the slope would be y2-y1/x2-x1 , right? YES

OpenStudy (beccaboo333):

still confused

OpenStudy (anonymous):

(3, 4) and (-3, -8) are your points (x1,y1) (x2,y2)

OpenStudy (anonymous):

and point slope form is (y - y1) = m (x-x1) you can find your m - the slope - using rise over run: the change in y divided by the change in x m = (y2 - y1)/(x2 - x1)

OpenStudy (beccaboo333):

uhm... what?

OpenStudy (beccaboo333):

how would i write the answer out?

OpenStudy (anonymous):

Your answer would be in the weird line form (y - y1)= m (x - x1) The y and the x that don't have subscripts are the variables that you normally see in an equation of a line. The y1 and x1 are the specific points the question gives you. (y - 4) = m (x - 3) Does that make more sense?

OpenStudy (beccaboo333):

is that both part one and two?

OpenStudy (beccaboo333):

Using complete sentences, explain whether or not it matters which point is used in the final answer. Also explain why you chose the point you did. thats part 2

OpenStudy (anonymous):

the slope m would be (y2 - y1)/(x2 - x1) = (-8 - 4)/(-3 -3) The second part is which point did you choose for the y1 and x1. It does matter which point you choose, because of how we defined the slope.

OpenStudy (beccaboo333):

so how would i write that?

OpenStudy (anonymous):

This all got very convoluted very quickly :P

OpenStudy (anonymous):

Well, does it make sense how the slope was defined?

OpenStudy (beccaboo333):

not at all. im not good with math..

OpenStudy (anonymous):

:) no worries. I'll draw a picture!

OpenStudy (beccaboo333):

but it has to be written

OpenStudy (anonymous):

I know, but it'll make it easier

OpenStudy (beccaboo333):

i need to get the assignment done soon or else itll cancel it :(

OpenStudy (anonymous):

harumph, fine. The slope is a number that shows how sharply the line is going up, so you use the change in y divided by the change in x, or (y2 - y1)(x2-x1)

OpenStudy (anonymous):

Since we also said that our equation of the line was (y - y1) = m (x - x1) we have to stay consistent because of how we defined m

OpenStudy (beccaboo333):

Thank you :) if i could give ya a hug for sticking around and helping i would.

OpenStudy (anonymous):

Glad this clusterbomb got sorted out :)

OpenStudy (beccaboo333):

thanks again

OpenStudy (anonymous):

good lucks and such

OpenStudy (beccaboo333):

thanks